# Measurement-induced transitions for interacting fermions

### v1

@@ I Introduction
An important question that has attracted considerable attention concerns the relation between charge fluctuations and information (entanglement). For random quantum circuits, it was argued that there are two distinct phase transitions that were termed “charge sharpening transition” and “purification transition”, respectively [27, 26, 30, 32] (see also [33]). Contrary to this, for free fermion models with measurements preserving the Gaussian character of a state, the charge and information are in direct relation (the second cumulant of particle number is proportional to the entanglement entropy), and thus exhibit the same behavior. It is, therefore, desirable to explore the measurement-induced phases and transitions between them in an interacting Hamiltonian model, for which a tractable analytical theory can be developed.
The goal of this paper is to understand the evolution of entanglement and charge fluctuations when the interaction is “switched on” in a free-fermion model with particle-number conservation. On the analytical side, we derive an NLSM field theory for interacting fermions and analyze it by renormalization-group (RG) means. This allows us to determine the phase diagram in the plane spanned by measurement rate and interaction. This phase diagram involves the transitions between an area-law phase and a phase with volume-law behavior for the entanglement entropy and between the area-law phase and the area×$\times$×log ($\ell^{d-1}\ln\ell$) phase for the particle-number cumulant. In particular, for the $d=1$ model, we demonstrate that interactions at not-too-frequent measurements destabilize the single area-law phase obtained previously for free fermions, giving rise to measurement-induced transitions in one dimension
@@ II.2 Observables
Both $S_{A}^{(N)}$ and $\chi_{A}^{(N)}(\lambda)$ are nonlinear functionals of the reduced density matrix, which makes averaging these objects over quantum trajectories nontrivial.
@@ III.1 Non-linear sigma model
is now indeed skew-symmetric: $\hat{H}_{0}^{\prime T}=-\hat{H}_{0}^{\prime}$, thus manifesting the PH symmetry. Such a transformation also works for any bipartite lattice with real hoppings; however, e.g., adding real next-nearest-neighbor hoppings (or non-uniform on-site energy) leads to the breakdown of the PH symmetry. We will be working with the transformed Hamiltonian $H_{0}^{\prime}$, dropping for brevity the prime symbol from now on.
@@ V.2 Berezinskii-Kosterlitz-Thouless transition for charge fluctuations in the symmetry-broken regime
V.3 Evolution of observables with increasing (sub)system size for 1D systems
Having performed the RG analysis of the interacting NLSM, we proceed with the predictions for the behavior of observables in finite systems (as realized in the numerical simulations). For definiteness, we will focus our discussion here on the case of 1D systems. Furthermore, since in most numerical simulations, the subsystem size is taken to be half the system size, $\ell_{A}=L/2$, we will consider the dependence of the entanglement entropy and the particle-number cumulant on $L$ for such half-cut chains. A generalization of our analysis on higher dimensions and other partitions is straightforward.
@@ V.3 Evolution of observables with increasing (sub)system size for 1D systems
We start with the smallest system sizes. Our predictions imply that both the charge fluctuations and entanglement in subsystems whose size is smaller or comparable to the “mean free path”, Eq. (64), $L\lesssim\ell_{0}$ (this condition is easily realized in not-too-large systems at weak monitoring), are governed by the ballistic dynamics. As was demonstrated in Ref. [53], the scaling of the particle-number cumulant and the entanglement entropy in a 1D free-fermion system in the ballistic regime is linear in $L$ (measured in units of the lattice constant):
$\begin{split}&\mathcal{S}\simeq-\left[n_{0}\ln n_{0}+(1-n_{0})\ln(1-n_{0})\right]\frac{L}{2},\quad L\lesssim\ell_{0},\\ &\mathcal{C}^{(2)}\simeq n_{0}(1-n_{0})\frac{L}{2},\quad L\lesssim\ell_{0}.\end{split}$ (125)
For half-filling, the prefactors in front of $L/2$ are equal to $\ln 2$ and $1/4$ for $\mathcal{S}$ and $\mathcal{C}^{(2)}$, respectively.
@@
Let us first consider the case $\ell_{\text{loc}}\ll\ell_{\text{int}}$. For a very high measurement rate, $\ell_{\text{loc}}\sim\ell_{0}$, the system gets immediately localized (the interaction does not play a role), showing the area-law behavior. For less frequent measurements, a window opens for the diffusive dynamics, $\ell_{0}\ll L\ll\ell_{\text{loc}}\ll\ell_{\text{int}}$, before localization sets in. The interaction-induced mass in replica-off-diagonal Goldstone modes can be neglected. The behavior of the charge fluctuations and entanglement follows the prediction [53] for non-interacting systems: the semiclassical logarithmic scaling is affected by the weak-localization correction (cf. Appendix C):
$\frac{3}{\pi^{2}}\,\mathcal{S}\simeq\mathcal{C}^{(2)}\simeq\frac{2}{\pi}\,g_{\text{WL}}(L)\ln L,\quad\ell_{0}\ll L\ll\ell_{\text{loc}},\ell_{\text{int}}\,,$ (127)
where $g_{\text{WL}}(L)$ is governed by Eq. (117). When $L$ reaches $\ell_{\text{loc}}$, a crossover to the area-law regime [Eq. (128)] occurs: both $\mathcal{C}^{(2)}$ and $\mathcal{S}$ do not depend on the system size (area law) when it exceeds $\ell_{\text{loc}}$:
$\mathcal{S},\,\mathcal{C}^{(2)}\sim g_{0}^{2}\propto\text{const}(L),\quad L\gtrsim\ell_{\text{loc}}\ \ (\ell_{\text{loc}}\ll\ell_{\text{int}}).$ (128)
@@
In the opposite limit, $\ell_{\text{int}}\ll\ell_{\text{loc}}$, the behavior of observables at scales between $\ell_{0}$ and $\ell_{\text{int}}$ is still barely affected by the interaction and is again described by Eq. (127). For $L\gtrsim\ell_{\text{int}}$, however, the situation changes drastically: the interaction now gaps some of the Goldstone modes and gives rise to charge-information separation. The coupling $g$ at $L\sim\ell_{\text{int}}$ is still large. The BKT RG, Eq. (122), then predicts that the renormalization of $g$ is negligible. The particle-number cumulant is described by the semiclassical result, Eq. (109), where $g$ is given by $g_{\text{WL}}(\ell_{\text{int}})$, corresponding to the purely logarithmic scaling of charge fluctuations in the thermodynamic limit:
$\mathcal{C}^{(2)}\simeq\frac{2}{\pi}\,g_{\text{WL}}(\ell_{\text{int}})\ln L,\quad L\gtrsim\ell_{\text{int}}\ \ (\ell_{\text{int}}\ll\ell_{\text{loc}}).$ (129)
Turning to the behavior of the entanglement entropy for $L\gtrsim\ell_{\text{int}}$ for the interaction-dominated regime $\ell_{\text{int}}\ll\ell_{\text{loc}}$, we use Eq. (111) with the renormalized value of $g$ at the length scale $\ell_{\text{int}}$. The asymptotic behavior of the entropy is then given by
$\mathcal{S}\simeq\frac{4}{3}g_{\text{WL}}(\ell_{\text{int}})\frac{L}{\ell_{\text{int}}},\quad L\gg\ell_{\text{int}}\ \ (\ell_{\text{int}}\ll\ell_{\text{loc}}),$ (130)
@@
Finally, we address the scale dependence of observables around $\ell_{\text{loc}}\sim\ell_{\text{int}}$, where the transition from metallic to insulating behavior takes place. In the semiclassical formula for the particle-number cumulant, Eq. (109), $g$ is now replaced with the decreasing coupling $g_{\text{BKT}}(L)$ governed by the BKT RG, Eq. (122). Exactly at the BKT separatrix, one has $g_{\text{BKT}}(\infty)=g_{c}=1/\pi$, yielding the critical scaling of the cumulant, Eq. (123). On the delocalized side of the charge-fluctuation transition, the coupling $g_{\text{BKT}}(L)$ reaches a constant value $g_{\text{BKT}}(\infty)>1/\pi$ (BKT line of fixed points) at the characteristic length $\ell_{\text{C}}$ given by Eq. (124). The scaling of the cumulant is described by Eq. (129), with a replacement $g_{\text{WL}}(\ell_{\text{int}})\mapsto g_{\text{BKT}}(\infty)$. When the initial condition for the BKT RG is on the opposite (“localized”) side of the separatrix, the charge fluctuations first follow the critical logarithmic scaling and then saturate at the scale $\ell_{\text{C}}$,
$\mathcal{C}^{(2)}\simeq\frac{2}{\pi^{2}}\ln\ell_{\text{C}},\quad g<g_{c},\ L\gg\ell_{\text{C}}\,.$ (131)
@@
As discussed in Sec. IV.3 above Eq. (111), the leading correction to the volume-law growth of $\mathcal{S}$ in the “metallic” regime, Eq. (130), is logarithmic in $L$. One can then speculate that the following behavior might persist near the transition: $\mathcal{S}\sim s_{L}\,L+s_{\text{log}}\ln L$, where $s_{L}$ and $s_{\text{log}}$ are coefficients depending on $\gamma$ and $V$, with $s_{L}$ vanishing at the transition $K=K_{c}$. Assuming that $s_{\text{log}}$ stays finite at the transition, the critical behavior of the entropy at the entanglement transition will be logarithmic, as in $d=1$ hybrid quantum circuits described by conformal field theories (see, e.g., Refs. [10, 22]). Near the transition for $K>K_{c}$, the volume-law asymptotics in the thermodynamic limit is established at $L\gtrsim\ell_{\text{E}}\gg\ell_{\text{int}}$. For stronger monitoring, $K<K_{c}$, the entropy growth is cut off at $\ell_{\text{E}}$ [Eq. (118)], leading to the crossover to the area law,
$\mathcal{S}=\text{const}(L),\quad K<K_{c},\ L\gg\ell_{\text{E}}.$ (132)
@@ V.4 Phase diagrams for 1D and 2D monitored interacting systems
As was discussed at the end of Sec. V.2, we cannot exclude a possibility that the two phase-transition lines actually coincide for small $V$ and $\gamma$ (where our derivation of the NLSM field theory holds). On the other hand, for $V\sim 1$ and $\gamma\sim 1$, the entanglement and charge fluctuations are decoupled already at the microscopic scale, making it very likely that the two transitions are distinct in this part of the phase diagram. There is thus a possibility that there is a single phase boundary for information and charge at small $V$ and $\gamma$, which then splits into two phase boundaries at a certain point where $V\sim 1$ and $\gamma\sim 1$.
@@ VI Numerical analysis
The main computational approach used in this work is a variant of the time-dependent variational principle (TDVP) for matrix product states (MPS) [129, 130], see Sec. VI.1, which has allowed us to proceed reliably up to system sizes $L=72$ in the relevant range of $\gamma$. As an auxiliary approach for studying density fluctuations, we also used time-dependent Hartree-Fock (TDHF) simulations [131], see Sec. VI.2.
@@
VI.1 Time-dependent variational principle with matrix product states (MPS-TDVP)
@@
Raising the bond dimension by a factor of two increases the maximum entropy by $\ln 2\approx 0.69$. This implies that the MPS calculation with $\chi_{\rm max}=256$, with $\mathcal{S}_{\rm max}(2^{8})\approx 5.54$, describes the system well up to $\mathcal{S}\approx 2.94$. In view of this, we limit the bond dimension to $\chi_{\max}=128$ for entropy values below $\mathcal{S}=2.25$ and use $\chi_{\max}=256$ in the range of parameters $(V,\gamma,L)$ where $\mathcal{S}>2.25$. Almost all our data satisfied the condition $\mathcal{S}<3$. For two points in the parameter space, where this condition was not fulfilled ($V=1$, $\gamma=0.3$, $L=54$ and 72), we carried out simulations with $\chi_{\max}=512$, which should give a good accuracy in the range $\mathcal{S}\leq 3.63$. This was well satisfied for $L=54$ but still slightly violated for $L=72$. We thus argue that all our TDVP-MPS results have good accuracy, with a somewhat higher numerical error expected for the latter point, see a detailed discussion below.
@@
To control the accuracy of our MPS-TDVP simulations, we used several complementary approaches. First, as mentioned above, we benchmarked the simulations for relatively small systems by comparing them to the ED results. Second, we performed a comparison to exact results obtained by means of time-evolved Green’s functions in the absence of interaction (cf. Appendix C), in which case the exact calculation can be carried out in the whole range of length used in our simulations [53]. Third, as also explained above, we verified the accuracy by performing simulations with different bond dimensions $\chi_{\max}$. Finally, to optimize our simulations, we used a variant of the TDVP algorithm that allows one to keep all bond dimensions variable throughout the simulation in an efficient way [133, 134, 135], see Appendix I for details.
@@
Let us first discuss the results for $V=0.5$. (dark blue curves). The highest considered measurement rate $\gamma=0.7$ is chosen such that the entropy and the particle-number cumulant can be seen to approach a plateau. The saturation of the entanglement entropy and the cumulant as a function of system size indicates that this point in the parameter space ($V=0.5$, $\gamma=0.7$) corresponds to the area-law phase. Furthermore, the (rescaled) cumulant $\tilde{\mathcal{C}}^{(2)}$ is almost identical to the entropy $\mathcal{S}$, as in the non-interacting case. This is exactly what is expected when the system is deeply in the area-law phase, with the “localization length” $\ell_{\text{loc}}$ being considerably smaller than the interaction-induced length $\ell_{\text{int}}$. The only visible effect of the interaction here is a relatively small ($\sim 8\%$) renormalization of the saturation value.
@@
VI.2 Time dependent Hartree-Fock (TDHF) simulatons

### v2

@@ I Introduction
An important question that has attracted considerable attention concerns the relation between charge fluctuations and information (entanglement). For random quantum circuits, it was argued that there are two distinct phase transitions that were termed “charge sharpening transition” and “purification transition”, respectively [27, 26, 30, 32] (see also [33, 113]). Contrary to this, for free fermion models with measurements preserving the Gaussian character of a state, the charge and information are in direct relation (the second cumulant of particle number is proportional to the entanglement entropy), and thus exhibit the same behavior. It is, therefore, desirable to explore the measurement-induced phases and transitions between them in an interacting Hamiltonian model, for which a tractable analytical theory can be developed.
Figure 1: Left panel: Schematic representation of a $d=1$ model, which includes nearest-neighbor hoppings (straight lines) and interactions (wavy lines), as well as a particle-number measurement apparatus attached to each site. Right panel: Phase diagram in the parameter plane of measurement rate $\gamma$ and interaction strength $V$, as obtained for $d=1$. From left to right: a phase with volume-law scaling for the entanglement entropy and logarithmic scaling for charge fluctuations (red), an intermediate phase with volume-law scaling for entanglement entropy and area-law scaling for charge fluctuations (magenta), and an area-law phase for both quantities (blue). Detailed description is provided in Sec. V.
The goal of this paper is to understand the evolution of entanglement and charge fluctuations when the interaction is “switched on” in a free-fermion model with particle-number conservation. Let us emphasize that charge fluctuations play a special role in models under consideration since the charge is the only conserved quantity. Indeed, both energy and momentum conservation laws are violated by quantum measurements. On the analytical side, we derive an NLSM field theory for interacting fermions and analyze it by RG means. This allows us to determine the phase diagram in the plane spanned by measurement rate and interaction. This phase diagram involves the transitions between an area-law phase and a phase with volume-law behavior for the entanglement entropy and between the area-law phase and the area×$\times$×log ($\ell^{d-1}\ln\ell$) phase for the particle-number cumulant. In particular, for the $d=1$ model, we demonstrate that interactions at not-too-frequent measurements destabilize the single area-law phase obtained previously for free fermions, giving rise to measurement-induced transitions in one dimension (Fig. 1).
@@ II.1 Generalized measurements
Formally, we will be interested in describing the ensemble in the steady-state regime defined by $T=t_{f}-t_{0}\to\infty$. It is worth noting that the limits $\gamma\to 0$ and $T\to\infty$ do not commute: the behavior of system at a weak measurement rate (small but finite $\gamma$) is drastically different from the behavior at $\gamma=0$ because, for any non-zero $\gamma$, the system evolves at a sufficiently long time $T$ to a measurement-induced steady state.
@@ II.2 Observables
Both $S_{A}^{(N)}$ and $\chi_{A}^{(N)}(\lambda)$ are nonlinear functionals of the reduced density matrix, which makes averaging these objects over quantum trajectories nontrivial. It is worth emphasizing that the ensemble-averaged density matrix has always a simple, infinite-temperature form and does not contain information about the observables of interest.
@@ III.1 Non-linear sigma model
is now indeed skew-symmetric: $\hat{H}_{0}^{\prime T}=-\hat{H}_{0}^{\prime}$, thus manifesting the PH symmetry. Such a transformation also works for any bi-partite lattice with real hoppings; however, e.g., adding real next-nearest-neighbor hoppings (or non-uniform on-site energy) leads to the breakdown of the PH symmetry. We will be working with the transformed Hamiltonian $H_{0}^{\prime}$, dropping for brevity the prime symbol from now on.
@@ V.2 Berezinskii-Kosterlitz-Thouless transition for charge fluctuations in the symmetry-broken regime
V.3 Evolution of observables with increasing subsystem size for 1D systems
Having performed the RG analysis of the interacting NLSM, we proceed with the predictions for the behavior of observables for finite subsystem size $\ell$, assuming thermodynamic limit for the size of the whole system $L\to\infty$. However, we note that the scaling with $\ell$ will be the same for the case $\ell=L/2$ (as realized in the numerical simulations) by virtue of continuity, although numerical prefactors may differ. For definiteness, we will focus our discussion here on the case of 1D systems. A generalization of our analysis on higher dimensions is straightforward.
@@ V.3 Evolution of observables with increasing subsystem size for 1D systems
We start with the smallest subsystem sizes. Our predictions imply that both the charge fluctuations and entanglement in subsystems whose size is smaller or comparable to the “mean free path”, Eq. (64), $\ell\lesssim\ell_{0}$ (this condition is easily realized in not-too-large systems at weak monitoring), are governed by the ballistic dynamics. As was demonstrated in Ref. [53], the scaling of the particle-number cumulant and the entanglement entropy in a 1D free-fermion system in the ballistic regime is linear in $\ell$ (measured in units of the lattice constant):
$\begin{split}&\mathcal{S}_{\ell}\simeq-\left[n_{0}\ln n_{0}+(1-n_{0})\ln(1-n_{0})\right]\ell,\quad\ell\lesssim\ell_{0},\\ &\mathcal{C}^{(2)}_{\ell}\simeq n_{0}(1-n_{0})\ell,\quad\ell\lesssim\ell_{0}.\end{split}$ (125)
For half-filling, the prefactors in front of $\ell$ are equal to $\ln 2$ and $1/4$ for $\mathcal{S}_{\ell}$ and $\mathcal{C}^{(2)}_{\ell}$, respectively.
@@
Let us first consider the case $\ell_{\text{loc}}\ll\ell_{\text{int}}$. For a very high measurement rate, $\ell_{\text{loc}}\sim\ell_{0}$, the system gets immediately localized (the interaction does not play a role), showing the area-law behavior. For less frequent measurements, a window opens for the diffusive dynamics, $\ell_{0}\ll\ell\ll\ell_{\text{loc}}\ll\ell_{\text{int}}$, before localization sets in. The interaction-induced mass in replica-off-diagonal Goldstone modes can be neglected. The behavior of the charge fluctuations and entanglement follows the prediction [53] for non-interacting systems: the semiclassical logarithmic scaling is affected by the weak-localization correction (cf. Appendix C):
$\frac{3}{\pi^{2}}\,\mathcal{S}_{\ell}\simeq\mathcal{C}^{(2)}_{\ell}\simeq\frac{2}{\pi}\,g_{\text{WL}}(\ell)\ln\ell,\quad\ell_{0}\ll\ell\ll\ell_{\text{loc}},\ell_{\text{int}}\,,$ (127)
where $g_{\text{WL}}(\ell)$ is governed by Eq. (117). When $\ell$ reaches $\ell_{\text{loc}}$, a crossover to the area-law regime [Eq. (128)] occurs: both $\mathcal{C}^{(2)}_{\ell}$ and $\mathcal{S}_{\ell}$ do not depend on the subsystem size (area law) when it exceeds $\ell_{\text{loc}}$:
$\mathcal{S}_{\ell},\,\mathcal{C}^{(2)}_{\ell}\sim g_{0}^{2}\propto\text{const}(\ell),\quad\ell\gtrsim\ell_{\text{loc}}\ \ (\ell_{\text{loc}}\ll\ell_{\text{int}}).$ (128)
@@
In the opposite limit, $\ell_{\text{int}}\ll\ell_{\text{loc}}$, the behavior of observables at scales between $\ell_{0}$ and $\ell_{\text{int}}$ is still barely affected by the interaction and is again described by Eq. (127). For $\ell\gtrsim\ell_{\text{int}}$, however, the situation changes drastically: the interaction now gaps some of the Goldstone modes and gives rise to charge-information separation. The coupling $g$ at $\ell\sim\ell_{\text{int}}$ is still large. The BKT RG, Eq. (122), then predicts that the renormalization of $g$ is negligible. The particle-number cumulant is described by the semiclassical result, Eq. (109), where $g$ is given by $g_{\text{WL}}(\ell_{\text{int}})$, corresponding to the purely logarithmic scaling of charge fluctuations in the thermodynamic limit:
$\mathcal{C}^{(2)}_{\ell}\simeq\frac{2}{\pi}\,g_{\text{WL}}(\ell_{\text{int}})\ln\ell,\quad\ell\gtrsim\ell_{\text{int}}\ \ (\ell_{\text{int}}\ll\ell_{\text{loc}}).$ (129)
Turning to the behavior of the entanglement entropy for $\ell\gtrsim\ell_{\text{int}}$ for the interaction-dominated regime $\ell_{\text{int}}\ll\ell_{\text{loc}}$, we use Eq. (111) with the renormalized value of $g$ at the length scale $\ell_{\text{int}}$. The asymptotic behavior of the entropy is then given by
$\mathcal{S}_{\ell}\simeq\frac{4}{3}g_{\text{WL}}(\ell_{\text{int}})\frac{\ell}{\ell_{\text{int}}},\quad\ell\gg\ell_{\text{int}}\ \ (\ell_{\text{int}}\ll\ell_{\text{loc}}),$ (130)
@@
Finally, we address the scale dependence of observables around $\ell_{\text{loc}}\sim\ell_{\text{int}}$, where the transition from metallic to insulating behavior takes place. In the semiclassical formula for the particle-number cumulant, Eq. (109), $g$ is now replaced with the decreasing coupling $g_{\text{BKT}}(\ell)$ governed by the BKT RG, Eq. (122). Exactly at the BKT separatrix, one has $g_{\text{BKT}}(\infty)=g_{c}=1/\pi$, yielding the critical scaling of the cumulant, Eq. (123). On the delocalized side of the charge-fluctuation transition, the coupling $g_{\text{BKT}}(\ell)$ reaches a constant value $g_{\text{BKT}}(\infty)>1/\pi$ (BKT line of fixed points) at the characteristic length $\ell_{\text{C}}$ given by Eq. (124). The scaling of the cumulant is described by Eq. (129), with a replacement $g_{\text{WL}}(\ell_{\text{int}})\mapsto g_{\text{BKT}}(\infty)$. When the initial condition for the BKT RG is on the opposite (“localized”) side of the separatrix, the charge fluctuations first follow the critical logarithmic scaling and then saturate at the scale $\ell_{\text{C}}$,
$\mathcal{C}^{(2)}_{\ell}\simeq\frac{2}{\pi^{2}}\ln\ell_{\text{C}},\quad g<g_{c},\ \ell\gg\ell_{\text{C}}\,.$ (131)
@@
As discussed in Sec. IV.3 above Eq. (111), the leading correction to the volume-law growth of $\mathcal{S}$ in the “metallic” regime, Eq. (130), is logarithmic in $\ell$. One can then speculate that the following behavior might persist near the transition: $\mathcal{S}_{\ell}\sim s_{\ell}\,\ell+s_{\text{log}}\ln\ell$, where $s_{\ell}$ and $s_{\text{log}}$ are coefficients depending on $\gamma$ and $V$, with $s_{\ell}$ vanishing at the transition $K=K_{c}$. Assuming that $s_{\text{log}}$ stays finite at the transition, the critical behavior of the entropy at the entanglement transition will be logarithmic, as in $d=1$ hybrid quantum circuits described by conformal field theories (see, e.g., Refs. [10, 22]). Near the transition for $K>K_{c}$, the volume-law asymptotics in the thermodynamic limit is established at $\ell\gtrsim\ell_{\text{E}}\gg\ell_{\text{int}}$. For stronger monitoring, $K<K_{c}$, the entropy growth is cut off at $\ell_{\text{E}}$ [Eq. (118)], leading to the crossover to the area law,
$\mathcal{S}_{\ell}=\text{const}(\ell),\quad K<K_{c},\ \ell\gg\ell_{\text{E}}.$ (132)
@@ V.4 Phase diagrams for 1D and 2D monitored interacting systems
As was discussed in the end of Sec. V.2, we cannot exclude a possibility that the two phase transition lines actually coincide for small $V$ and $\gamma$ (where our derivation of the NLSM field theory holds). On the other hand, for $V\sim 1$ and $\gamma\sim 1$, the entanglement and charge fluctuations are decoupled already at the microscopic scale, making it very likely that the two transitions are distinct in this part of the phase diagram. There is thus a possibility that there is a single phase boundary for information and charge at small $V$ and $\gamma$, which then splits into two phase boundaries at a certain point where $V\sim 1$ and $\gamma\sim 1$.
@@ VI Numerical analysis
The main computational approach used in this work is a variant of the time-dependent variational principle for matrix product states [130, 131], see Sec. VI.1, which has allowed us to proceed reliably up to system sizes $L=72$ in the relevant range of $\gamma$. As an auxiliary approach for studying density fluctuations, we also used time-dependent Hartree-Fock simulations [132], see Sec. VI.2.
@@
VI.1 Time-dependent variational principle with matrix product states
@@
Raising the bond dimension by a factor of two increases the maximum entropy by $\ln 2\approx 0.69$. This implies that the MPS calculation with $\chi_{\rm max}=256$, with $\mathcal{S}_{\rm max}(2^{8})\approx 5.54$, describes the system well up to $\mathcal{S}\approx 2.94$. In view of this, we limit the bond dimension to $\chi_{\max}=128$ for entropy values below $\mathcal{S}=2.25$ and use $\chi_{\max}=256$ in the range of parameters $(V,\gamma,L)$ where $\mathcal{S}>2.25$. Almost all our data satisfied the condition $\mathcal{S}<3$. For two points in the parameter space, where this condition was not fulfilled ($V=1$, $\gamma=0.3$, $L=54$ and 72), we carried out simulations with $\chi_{\max}=512$, which should give a good accuracy in the range $\mathcal{S}\leq 3.63$. This was well satisfied for $L=54$ but still slightly violated for $L=72$. We thus argue that all our MPS-TDVP results have good accuracy, with a somewhat higher numerical error expected for the latter point, see a detailed discussion below.
@@
To control the accuracy of our MPS-TDVP simulations, we used several complementary approaches. First, as mentioned above, we benchmarked the simulations for relatively small systems by comparing them to the ED results. Second, we performed a comparison to exact results obtained by means of time-evolved Green’s functions in the absence of interaction (cf. Appendix C), in which case the exact calculation can be carried out in the whole range of length used in our simulations [53]. Third, as also explained above, we verified the accuracy by performing simulations with different bond dimensions $\chi_{\max}$. The saturation with respect to the bond dimension is illustrated in Fig. 17 of Appendix I for the case of the largest entanglement in our simulations (largest system size, $L=72$, strongest interaction, $V=1$, and smallest $\gamma=0.3$). Finally, to optimize our simulations, we used a variant of the TDVP algorithm that allows one to keep all bond dimensions variable throughout the simulation in an efficient way [134, 135, 136], see Appendix I for details.
@@
Let us first discuss the results for $V=0.5$ (dark blue curves). The highest considered measurement rate $\gamma=0.7$ is chosen such that the entropy and the particle-number cumulant can be seen to approach a plateau. The saturation of the entanglement entropy and the cumulant as a function of system size indicates that this point in the parameter space ($V=0.5$, $\gamma=0.7$) corresponds to the area-law phase. Furthermore, the (rescaled) cumulant $\tilde{\mathcal{C}}^{(2)}$ is almost identical to the entropy $\mathcal{S}$, as in the non-interacting case. This is exactly what is expected when the system is deeply in the area-law phase, with the “localization length” $\ell_{\text{loc}}$ being considerably smaller than the interaction-induced length $\ell_{\text{int}}$. The only visible effect of the interaction here is a relatively small ($\sim 8\%$) renormalization of the saturation value.
@@
VI.2 Time dependent Hartree-Fock simulations